OpenStudy (anonymous):

Simplifying radical expressions: \[\frac{\sqrt(2) + \sqrt(5)}{\sqrt(2) - \sqrt(5)}\] Here is my attempt: \[\frac{\sqrt(2) + \sqrt(5) * \sqrt(2) - \sqrt(5)}{\sqrt(2) - \sqrt(5) * \sqrt(2) - \sqrt(5)}\] (Multiplying via the conjugate.) I am not sure where to go from there, the question asks me to simplify to the simplest form. Any help is appreciate, Thanks.

OpenStudy (anonymous):

you need to multiply by the conjugate of the denominator, not the numerator

OpenStudy (anonymous):

\[\frac{\sqrt2 + \sqrt5}{\sqrt2 - \sqrt5}\frac{\sqrt2+\sqrt5}{\sqrt2+\sqrt5}\] is the first step

OpenStudy (anonymous):

the denominator is \(2-5=-3\) and the numerator is what you get when you multiply that mess out

OpenStudy (anonymous):

\[\dfrac {7 + 2 \sqrt(10)}{-3}\] My work: \[{(\sqrt(2) + \sqrt(5))}^2 = \sqrt(2)^2 + 2\sqrt(2)\sqrt(5) + \sqrt(5)^2\] \[7 + 2 \sqrt(10)\] \[2 - 5 = -3\] \[\dfrac{7 + 2 \sqrt(10)}{-3}\] Thank you, I understand. Brilliant!

OpenStudy (anonymous):

looks good to me